I need to graph the quadratic function f(x)= -1/2(x+4)^2+6
I've determined the parabola opens down due to the negative coefficient; however, I'm having a problem solving it due to the fraction -1/2. Wouldn't (x+4)^2 be x^2+16?

Answer by jim_thompson5910(35256)

(x+4)^2 =(x+4)(x+4) = x*x + 4x + 4x + 4*4 = x^2 + 4x + 4x + 16 = x^2 + 8x + 16


So (x+4)^2 = x^2 + 8x + 16 for all values of x.